40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 3 21:05:23 2022, yuta, Update, BHD, AS path alignment, removing a lot of green stuff 7x
    Reply  Wed May 4 13:09:42 2022, Anchal, Update, BHD, IMC table balanced again, IMC is locking, YARM is locking 
       Reply  Wed May 4 15:08:59 2022, Tega, Update, BHD, XARM alignment to get flashing IFO_XY_ARMS_Flashing.png
          Reply  Wed May 4 18:52:43 2022, Tega, Update, BHD, POP beam too high at POP SM4 on the ITMX chamber snnp_pr2_pr3_itmy_aligment.png
             Reply  Thu May 5 14:46:22 2022, Anchal, Update, BHD, POP beam height lowered, POP_SM4 raised 
                Reply  Thu May 5 17:05:31 2022, Tega, Update, BHD, IMC & X/Y-arm alignment IFO_aligment_Y_locking_on.png
                Reply  Mon May 9 18:43:03 2022, Anchal, Update, BHD, ITMX table layout corrected 
                   Reply  Mon May 9 18:49:05 2022, Tega, Update, BHD, Relocate green TRX and TRY PDs/cams/optics from PSL table to BS table IMG_20220509_184439943.jpgIMG_20220509_184154722.jpg
                      Reply  Tue May 10 15:46:38 2022, JC, Update, BHD, Relocate green TRX and TRY components from PSL table to BS table IMG_0741.jpegIMG_0753.jpeg
                         Reply  Tue May 10 18:25:43 2022, Paco, Update, BHD, Green Transmission path 
                   Reply  Tue May 10 16:31:14 2022, Tega, Update, BHD, ITMX optlev return beam steered to QPD 
       Reply  Fri May 6 13:48:34 2022, Anchal, Update, BHD, WFS loop instability fixed 
    Reply  Thu May 12 19:55:01 2022, Tega, Update, BHD, AS path alignment AS.JPGOMCchamber.JPG
       Reply  Thu May 12 20:11:18 2022, Anchal, Update, BHD, BHDBS Output beams steered out to ITMY table signal-2022-05-12-201844.jpeg
          Reply  Thu May 12 20:24:29 2022, yuta, Update, BHD, POX and POY investigation POY.JPG
       Reply  Mon May 16 13:22:59 2022, yuta, Update, BHD, REFL and AS paths aligned at AP table REFLPOP.JPGPOPAS.JPG
          Reply  Mon May 16 19:14:17 2022, Anchal, Update, BHD, Camera set on AS path and BHDBS output path 
             Reply  Tue May 17 18:43:38 2022, Anchal, Update, BHD, Placed SRM in ITMY Chamber 
                Reply  Wed May 18 08:30:29 2022, Paco, Update, BHD, SRM OpLev 
                   Reply  Thu May 19 08:51:40 2022, Paco, Update, BHD, SRM OpLev 
                Reply  Wed May 18 09:02:52 2022, Anchal, Update, BHD, WFS1 PD centered 
                Reply  Wed May 18 17:23:15 2022, Anchal, Update, BHD, Placed PRM in BS Chamber 
                   Reply  Thu May 19 18:56:08 2022, yuta, Update, BHD, Current OSEM sensor values with all the suspensions aligned OSEMS.txt
                   Reply  Thu May 19 19:05:59 2022, Anchal, Update, BHD, BS Chamber all work finished, BHD path setup 
                      Reply  Fri May 20 20:03:48 2022, Paco, Update, BHD, ITMY chamber work finished - LO and AS overlapped Screenshot_2022-05-20_20-38-25.pngOSEMs.txt
                         Reply  Mon May 23 13:16:59 2022, Paco, Update, BHD, End of vent - checks OSEMs20220523.txtScreenshot_2022-05-23_13-24-06_IFOaligned_GTRY_still_clipping_but_less.pngPR3position.JPG
                            Reply  Tue May 24 15:21:13 2022, Anchal, Update, BHD, Freeswing tests of new SOS started 
Message ID: 16831     Entry time: Wed May 4 18:52:43 2022     In reply to: 16830     Reply to this: 16832
Author: Tega 
Type: Update 
Category: BHD 
Subject: POP beam too high at POP SM4 on the ITMX chamber 

[Yuta, Tega]

We needed to sort out the POXDC signal so we could work on X-arm alignment. Given that POXDC channel value was approx 6 compared to POYDC value of approx. 180, we decided to open the ITMX chamber to see if we could improve the situation. We worked on the alignment of POX beam but could not improve the DC level which suggests that this was already optimized for.  As an aside, we also noticed some stray IR beam from the BS chamber, just above the POX beam which we cold not identify.

Next we moved on to the POP beam alignment, where we noticed that the beam level on LO1 and POP_SM4 was a bit on the high side. Basically, the beam was completely missing the 1" POP_SM4 mirror and was close to the top edge of LO1. So we changed TT2 pitch value from 0.0143 to -0.2357 in order to move the beam position on POP_SM4 mirror. This changed the input alignment, so we compensated using PR2 (0.0 -> 49.0) and PR3 (-5976.560 -> -5689.800). This did not get back the alignment as anticipated, so we moved ITMY pitch from 0.9297 to 0.9107. All of these alignment changes moved the POP beam down by approx 1/5 of an inch from outside the mirro to the edge of POP_SM4 mirror, where about half of the beam is clipped.

 

Next Steps:

We need to repeat these aligment procedures with say 1.5 time the change in TT2 pitch to center the beam on POP_SM4 mirror.

Attachment 1: snnp_pr2_pr3_itmy_aligment.png  167 kB  Uploaded Wed May 4 20:06:29 2022  | Hide | Hide all
snnp_pr2_pr3_itmy_aligment.png
ELOG V3.1.3-